OpenStudy (anonymous):

solve system of equation: x+y=-1 y-z=4 2x+y+5z=-38 i need step by step help please?

OpenStudy (campbell_st):

ok... take the 2nd equation y - z = 4 and multiply it by 5 so you get 5y - 5z = 20 now add it to the last equation this will eliminate z 5y - 5z = 20 2x + y + 5z = -38 ------------------ 2x + 6y = -18 now you have 2 equations in 2 unknowns... which is easier to solve you can divide the equation by 2 to make easier x + 3y = -9 and x + y = -1 a little subtraction x + 3y = -9 -x - y = 1 ------------- 2y = -8 so y = -4 substitute the value of y into the 1st equation to find x x - 4 = -1 when you have x substitute your value of y into the 2nd equation to find z -4 - z = 4 hope this helps

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

and z is equal t0 7 corretc?

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

no I think z = -8

OpenStudy (anonymous):

and x is 3 correct?

OpenStudy (campbell_st):

yep... I think thats it

OpenStudy (anonymous):

thank yo so much. You work fast :)